OpenStudy (anonymous):

Inverse the function f(x) = 2x/x-1 I have so far: y = 2x/x-1 y(x-1) = 2x But I'm stuck here, I know I need to solve for x, but I don't see what I can do to achieve it.

OpenStudy (shamim):

ok i wanna try to solve it

OpenStudy (shamim):

\[y=\frac{ 2x }{ x-1 }\]

OpenStudy (shamim):

\[or, y(x-1)=2x\]

OpenStudy (shamim):

\[or, xy-y=2x\]

OpenStudy (shamim):

r u following me

OpenStudy (shamim):

\[or, xy-2x=y\]

OpenStudy (shamim):

\[x(y-2)=y\]

OpenStudy (shamim):

\[or, x=\frac{ y }{ y-2 }\]

OpenStudy (shamim):

is it ok

OpenStudy (shamim):

plz tell me somethig

OpenStudy (shamim):

f inverse y=x=y/(y-2)

OpenStudy (anonymous):

Yeah, that works. I completely blanked on adding/subtracting from the equation